I am not sure how to solve this problem. I'd appreciate the help, thank you!!
"A scoring possibility in a game of Yahtzee with five 6-sided dice is called a "three of a kind," where three numbers are the same and the remaining two numbers are both different (eg: 5 5 5 1 4 or 2 5 2 3 2). Find the probability of rolling a three of a kind.

Answer by stanbon(75887)

"A scoring possibility in a game of Yahtzee with five 6-sided dice is called a "three of a kind," where three numbers are the same and the remaining two numbers are both different (eg: 5 5 5 1 4 or 2 5 2 3 2). Find the probability of rolling a three of a kind.
---
Pick a number: 6 ways
Pick 3 of that number on 5 dice: 5C3 = 10 ways
Pick another number: 5 ways
Pick yet another number: 4 ways
Total # of ways to succeed: 6*10*5*4 = 1200
---
# of possible outcomes: 6^5 = 7776
---
Probability of "three of a kind" = 1200/7776 = 0.1543
=======================================
Cheers,
Stan H.